# A Probabilistic Approach to Uncertainty Quantification Leveraging 3D Geometry

### v1

@@ Abstract
Quantifying uncertainty in neural implicit 3D representations, particularly those utilizing Signed Distance Functions (SDFs), remains a substantial challenge due to computational inefficiencies, scalability issues, and geometric inconsistencies. Existing methods typically neglect direct geometric integration, leading to poorly calibrated uncertainty maps. We introduce BayesSDF, a novel probabilistic framework for uncertainty quantification in neural implicit SDF models, motivated by scientific simulation applications with 3D environments (e.g., forests) such as modeling fluid flow through forests, where precise surface geometry and awareness of fidelity surface geometric uncertainty are essential. Unlike radiance-based models such as NeRF or 3D Gaussian splatting, which lack explicit surface formulations, SDFs define continuous and differentiable geometry, making them better suited for physical modeling and analysis. BayesSDF leverages a Laplace approximation to quantify local surface instability via Hessian-based metrics, enabling computationally efficient, surface-aware uncertainty estimation. Our method shows that uncertainty predictions correspond closely with poorly reconstructed geometry, providing actionable confidence measures for downstream use. Extensive evaluations on synthetic and real-world datasets demonstrate that BayesSDF outperforms existing methods in both calibration and geometric consistency, establishing a strong foundation for uncertainty-aware 3D scene reconstruction, simulation, and robotic decision-making.
@@ 1 Introduction
Uncertainty quantification is a fundamental challenge in machine learning, particularly in 3D scene reconstruction, where ensuring reliability in learned representations is crucial for downstream applications such as robotics, medical imaging, and physical simulations. Neural implicit representations have revolutionized 3D geometry learning by encoding scenes as continuous functions rather than discrete point clouds or meshes. Among these, Neural Radiance Fields (NeRF) [5] enable photorealistic novel view synthesis but lack explicit geometric constraints, making them less suitable for precise surface reconstruction. To address this limitation, Neural Implicit Surfaces (NeuS) [7] employ Signed Distance Functions (SDFs), providing a more structured representation of 3D geometry by modeling surfaces as zero-level sets of a continuous field. This geometric formulation allows NeuS to achieve superior surface consistency and geometric regularity, but the challenge remains in effectively quantifying uncertainty in the learned SDF field. In this paper, we introduce BayesSDF, a novel probabilistic framework that enables surface-aligned uncertainty estimation in SDF-based 3D neural representations. Rather than relying on ensembles or pixel-level sampling, BayesSDF utilizes a Laplace approximation over a learned deformation field to compute a Hessian-based metric that captures the local sensitivity of the rendering function. This formulation ensures that uncertainty maps are geometrically coherent. By integrating the geometric advantages of SDF-enabled models with a principled probabilistic framework, BayesSDF establishes an unexplored view for uncertainty-aware 3D scene reconstruction and representation learning. The method enables precise identification of areas with high reconstruction variance, leading to better surface-aware uncertainty estimates. BayesSDF also improves calibration and geometric consistency over previous methods. Instead of estimating uncertainty through pure pixel-space renderings or depth-based sampling like NeRF, BayesSDF utilizes the continuous SDF field from geometrically adaptable reconstruction models, reducing computational overhead while maintaining accuracy. The proposed method demonstrates superior performance across multiple synthetic and real-world datasets, achieving state-of-the-art uncertainty calibration and reliability in 3D surface modeling.
@@ 2 Contributions
This research introduces significant contributions to the field of uncertainty quantification for 3D implicit neural representations:
@@
    Direct Geometric Integration: Unlike existing techniques that indirectly approximate uncertainty through pixel-space or depth-based approaches, BayesSDF explicitly integrates the geometric properties inherent in Signed Distance Functions. This direct integration significantly improves geometric consistency and calibration of uncertainty predictions.
@@
    Efficient Hessian-based Uncertainty Estimation: BayesSDF employs a Hessian-based uncertainty calculation method that efficiently captures local sensitivity in the geometry deformation field. By evaluating how small perturbations in geometric parameters affect rendered color outputs, the method accurately quantifies spatial uncertainty without the computational expense of traditional ensemble or extensive sampling approaches.
@@
    Comprehensive Evaluation and Validation: Extensive empirical validation on synthetic and real-world datasets demonstrates BayesSDF’s superiority over existing uncertainty quantification methods. Metrics including the Area Under Sparsification Error (AUSE) show that BayesSDF achieves comparable or superior calibration performance, especially notable in complex real-world scenarios.
@@
    Scientific Motivation for SDF-Based Uncertainty: We motivate our approach by targeting applications in scientific simulation and computational physics, where accurate geometry and interpretable uncertainty are critical. SDF-based models, as demonstrated in DeepSDF [6], provide surface continuity and differentiability that are advantageous for physical modeling. Our framework explicitly supports such use cases by tying uncertainty to local surface instability.
The outlined contributions solidify BayesSDF as a robust, geometrically coherent, and computationally efficient uncertainty quantification framework with significant implications across various domains such as scene reconstruction, physics-based simulations, and autonomous robotic navigation.
@@ 3 Related Work
Uncertainty quantification in neural implicit representations is a rapidly evolving field. Traditional deep learning models often produce deterministic outputs without explicit confidence measures, making them unreliable in critical scenarios. In the context of 3D learning, uncertainty quantification plays a crucial role in assessing reconstruction accuracy, identifying ambiguous regions, and guiding active learning. Below, we review the most relevant areas of uncertainty quantification in 3D learning, including ensemble-based methods, neural radiance fields (NeRF), neural implicit surfaces (NeuS), signed distance functions (SDFs), Bayesian uncertainty estimation techniques, and more. We then highlight how BayesSDF significantly advances the field by integrating uncertainty directly into SDF-based neural implicit models, enabling computationally efficient and geometrically consistent uncertainty estimation.
@@ 3.1 Uncertainty Estimation in 3D Reconstruction Methods
Uncertainty estimation plays a crucial role in 3D reconstruction, where accurate surface modeling is essential for downstream applications. Various methods exist for reconstructing 3D surfaces and each of these approaches presents unique challenges for uncertainty quantification.
@@ 3.1.1 Neural Representations and SDF-Based Methods
Neural implicit representations, such as NeRF (Neural Radiance Fields) and NeuS (Neural Implicit Surfaces), have demonstrated state-of-the-art performance in view synthesis and geometric reconstruction. However, their uncertainty estimation techniques differ significantly:
@@
  • NeuS-based methods instead represent surfaces using a signed distance function (SDF), which provides a continuous and differentiable description of 3D geometry. However, existing uncertainty estimation techniques in NeRF do not translate well to NeuS, as they rely on ray-space density variations rather than surface-level geometric uncertainty.
@@ 3.2 Uncertainty Estimation in Ensemble-Based Learning
  • Computational Cost: Training multiple neural implicit models is memory and compute intensive, particularly for high-resolution 3D reconstructions. Traditional ensemble methods require storing and evaluating multiple models, leading to scalability issues.
@@
  • Ray vs. Volume-Based Estimation: Many existing ensemble-based uncertainty estimation techniques in NeRF-like models operate strictly along rays. This results in uncertainty that does not align well with actual geometric structures, making it less interpretable.
By leveraging the strengths of ensemble-based learning while addressing key limitations in 3D implicit representations, BayesSDF establishes a new state-of-the-art paradigm for uncertainty quantification in neural implicit fields.
@@ 3.3 Bayesian Inference For Ray-Based sampling
BayesRays [2] is a recent approach for uncertainty quantification in neural radiance fields. It modifies NeRF by introducing a Bayesian framework that estimates uncertainty in the color and density outputs of the model. While BayesRays is effective for quantifying NeRF uncertainty, it remains tied to volumetric density representations and does not explicitly model surface uncertainty. Additionally, BayesRays relies on ray-space sparsification techniques, which can lead to misalignment between estimated uncertainty and actual geometric uncertainty. BayesSDF rebuilds this previously established ray sampling uncertainty method by integrating reconstruction methods that inherently reveal volumetric qualities to ensure direct alignment between uncertainty estimates and the scene geometry, putting less emphasis on the reliance of rendering estimations to make it inherently more interpretable for geometric reconstruction tasks, as it latently exploits surface variability rather than depth-based uncertainty.
3.4 Comparison of Existing Methods
@@
Table 1: Comparison of Existing 3D Reconstruction Methods in Terms of Representation, Uncertainty Estimation, Computational Cost, and Geometric Consistency.
@@ 5.1 Theory
Our goal is to model an associated uncertainty measure, derived from the local sensitivity of the rendering function to small perturbations in the underlying deformation parameters. Let $\mathbf{x}\in\mathbb{R}^{3}$ denote a 3D point in scene space, and let $\mathbf{c}(\mathbf{x})\in\mathbb{R}^{3}$ be the color (in RGB) rendered by the neural field. Rather than relying purely on a deterministic mapping $\mathbf{x}\mapsto\mathbf{c}(\mathbf{x})$, we attach a set of local deformation parameters $\mathbf{d}(\mathbf{x})$ that influence the geometry (and thus the color) through a learned distortion or deformation field. Formally, we assume $\mathbf{c}(\mathbf{x})\;=\;\mathcal{R}\bigl{(}\mathbf{x},\mathbf{d}(\mathbf{x});\Theta\bigr{)}$ where $\mathcal{R}$ is the rendering operator that depends on both the scene geometry (parameterized by a neural SDF model and any additional shape parameters $\Theta)$ and the local deformation $\mathbf{d}(\mathbf{x})$. In order to estimate uncertainty, we measure how sensitively $\mathbf{c}(\mathbf{x})$ changes with respect to changes in $\mathbf{d}(\mathbf{x})$. Intuitively, if small variations in $\mathbf{d}(\mathbf{x})$ induce large changes in color, we interpret that region as having higher uncertainty in the final scene reconstruction.
5.2 Leveraging Distortion Fields
@@
The distortion field $f_{\mathrm{hash}}(\mathbf{x})$ captures spatial distortions in the SDF field, serving as a basis for uncertainty estimation. To represent the deformation parameters $\mathbf{d}(\mathbf{x})$ on a continuous domain, we employ a HashEncoding to serve as a voxel-like grid but implemented through a hash table $\mathbf{d}(\mathbf{x})=f_{\mathrm{hash}}(\mathbf{x})\in\mathbb{R}^{3}$ where $f_{\mathrm{hash}}$ is a function learned by a single-level hash encoding with: $\mathrm{min\_res}=\mathrm{max\_res}=2^{\ell}$ for level-of-detail $\ell$, a hash table size of $2^{3\ell+1}$ entries, a linear interpolation scheme for values between grid corners, and 3 output channels per queried grid vertex. Effectively, for a given point $\mathbf{x}$, the hash encoding looks up an 8-corner neighborhood around the voxel containing $\mathbf{x}$, interpolates among these corners (each storing a 3D deformation vector), and outputs a single 3D vector $\mathbf{d}(\mathbf{x})$. This approach is parameter-efficient while remaining flexible. We denote $\mathbf{aabb}\in\mathbb{R}^{6}$ as the scene bounding box, which is used to normalize $\mathbf{x}$ to a unit cube or appropriate domain before querying the hash grid. Specifically, we generate a ray bundle from a dataset’s rays, which contains origins, directions, and near/far bounds. We then use Neus’ hierarchical or “proposal” sampler to refine sample locations along each ray, focusing more samples near scene geometry (i.e., near the SDF zero level set). Let $\{\mathbf{x}_{i}\}$ be the set of sampled 3D points on a ray. When each ray sample $\mathbf{x}\in\mathbb{R}^{3}$ is generated, we transform $\mathbf{x}$ into normalized coordinates $\tilde{\mathbf{x}}=\mathrm{normalize}(\mathbf{x},\mathbf{aabb})$. Then, $\mathbf{d}(\mathbf{x})\;=\;\sum_{c=1}^{8}\alpha_{c}(\tilde{\mathbf{x}})\,\mathbf{d}_{c}\quad\text{where}\quad\bigl{\{}\mathbf{d}c\bigr{\}}{c=1}^{8}\;$ are are the corner deformation vectors that are found from a grid cell index in each dimension (one for each binary combination of $\lfloor\cdot\rfloor or\lceil\cdot\rceil$ in (x,y,z)) and $\alpha_{c}(\tilde{\mathbf{x}})$ are the trilinear interpolation coefficients $(\alpha_{c}(\mathbf{x}i)\;\in\;[0,1],\quad c=1,\ldots,8,)$ such that $\sum{c=1}^{8}\alpha_{c}(\mathbf{x}_{i})=1$. These coefficients are used to blend the deformation vectors from the 8 corners. This procedure is vital not only for generating $\mathbf{d}(\mathbf{x}_{i})$, but also for accumulating partial derivatives. Because each point’s deformation influences eight corner entries, we must carefully track which corners are involved and in what proportion.

### v2

# BayesSDF: Surface-Based Laplacian Uncertainty Estimation for 3D Geometry with Neural Signed Distance Fields

@@ Abstract
Quantifying uncertainty in neural implicit 3D representations, particularly those utilizing Signed Distance Functions (SDFs), remains a substantial challenge due to computational inefficiencies, scalability issues, and geometric inconsistencies. Existing methods typically neglect direct geometric integration, leading to poorly calibrated uncertainty maps. We introduce BayesSDF, a novel probabilistic framework for uncertainty quantification in neural implicit SDF models, motivated by scientific simulation applications with 3D environments (e.g., forests) such as modeling fluid flow through forests, where precise surface geometry and reliable uncertainty estimates are essential. Unlike radiance-based models such as Neural Radiance Fields (NeRF) or 3D Gaussian splatting, which lack explicit surface formulations, Signed Distance Functions (SDFs) define continuous and differentiable geometry, making them better suited for physical modeling and analysis. BayesSDF leverages a Laplace approximation to quantify local surface instability using Hessian-based metrics, enabling efficient, surfaceaware uncertainty estimation. Our method shows that uncertainty predictions correspond closely with poorly reconstructed geometry, providing actionable confidence measures for downstream use. Extensive evaluations on synthetic and real-world datasets demonstrate that BayesSDF outperforms existing methods in both calibration and geometric consistency, establishing a strong foundation for uncertainty-aware 3D scene reconstruction, simulation, and robotic decision-making.
@@ 1 Introduction
Uncertainty quantification is a fundamental challenge in machine learning, particularly in 3D scene reconstruction, where ensuring reliability in learned representations is crucial for downstream applications such as robotics, medical imaging, and physical simulations. Neural implicit representations have revolutionized 3D geometry learning by encoding scenes as continuous functions rather than discrete point clouds or meshes. Among these, Neural Radiance Fields (NeRF) [5] enable photorealistic novel view synthesis but lack explicit geometric constraints, making them less suitable for precise surface reconstruction. To address this limitation, Neural Implicit Surfaces (NeuS) [7] employ Signed Distance Functions (SDFs), providing a more structured representation of 3D geometry by modeling surfaces as zero-level sets of a continuous field. This geometric formulation allows NeuS to achieve superior surface consistency and geometric regularity, but the challenge remains in effectively quantifying uncertainty in the learned SDF field. In this paper, we introduce BayesSDF, a novel probabilistic framework that enables surface-aligned uncertainty estimation in SDF-based 3D neural representations. Rather than relying on ensembles or pixel-level sampling, BayesSDF utilizes a Laplace approximation over a learned deformation field to compute a Hessian-based metric that captures the local sensitivity of the rendering function. This formulation ensures that uncertainty maps are geometrically coherent. By integrating the geometric advantages of SDF-enabled models with a principled probabilistic framework, BayesSDF establishes an unexplored view for uncertainty-aware 3D scene reconstruction and representation learning. The method enables precise identification of areas with high reconstruction variance, leading to improved surface-aware uncertainty estimates. BayesSDF also improves calibration and geometric consistency over previous methods. Instead of estimating uncertainty through pure pixel-space renderings or depth-based sampling like NeRF, BayesSDF utilizes the continuous SDF field from geometrically adaptable reconstruction models, reducing computational overhead while maintaining accuracy. The proposed method demonstrates superior performance across multiple synthetic and real-world datasets, achieving state-of-the-art uncertainty calibration and reliability in 3D surface modeling. This research introduces significant contributions to the field of uncertainty quantification for 3D implicit neural representations:
@@ 2 Contributions
    Direct Geometric Integration: Unlike existing techniques that indirectly approximate uncertainty through pixel-space or depth-based approaches, BayesSDF explicitly integrates the geometric properties inherent in Signed Distance Functions. This direct integration significantly improves the geometric consistency and calibration of uncertainty predictions.
@@
    Efficient Hessian-based Uncertainty Estimation: BayesSDF employs a Hessian-based uncertainty calculation method that efficiently captures local sensitivity in the geometry deformation field. By evaluating how small perturbations in geometric parameters affect rendered color outputs, the method accurately quantifies spatial uncertainty, avoiding the computational expense of traditional ensemble or extensive sampling approaches.
@@
    Comprehensive Evaluation and Validation: Extensive empirical validation on synthetic and real-world datasets demonstrates BayesSDF’s superiority over existing uncertainty quantification methods. Metrics such as the Area Under Sparsification Error (AUSE) show that BayesSDF achieves comparable or superior calibration performance, especially in complex real-world scenarios.
@@
    Scientific Motivation for SDF-Based Uncertainty: We motivate our approach by targeting applications in scientific simulation and computational physics, where accurate geometry and interpretable uncertainty are critical. SDF-based models, as demonstrated in DeepSDF [6], provide surface continuity and differentiability that are advantageous for physical modeling. Our framework explicitly supports such use cases by relating uncertainty to local surface instability.
These contributions establish BayesSDF as a robust, geometrically coherent, and computationally efficient uncertainty quantification framework with significant implications for scene reconstruction, physics-based simulations, and autonomous robotic navigation.
@@ 3 Related Work
Uncertainty quantification in neural implicit representations is a rapidly evolving field. Traditional deep learning models often produce deterministic outputs without explicit confidence measures, making them unreliable in critical scenarios. In the context of 3D learning, uncertainty quantification plays a crucial role in assessing reconstruction accuracy, identifying ambiguous regions, and guiding active learning. Below, we review the most relevant areas of uncertainty quantification in 3D learning, including ensemble-based methods, neural radiance fields (NeRF), neural implicit surfaces (NeuS), signed distance functions (SDFs), Bayesian uncertainty estimation techniques, among others. We then highlight how BayesSDF significantly advances the field by integrating uncertainty directly into SDF-based neural implicit models, enabling computationally efficient and geometrically consistent uncertainty estimation.
@@ 3.1 Uncertainty Estimation in 3D Reconstruction Methods
Uncertainty estimation plays a crucial role in 3D reconstruction, where accurate surface modeling is essential for downstream applications. Various methods exist for reconstructing 3D surfaces, and each presents unique challenges for uncertainty quantification.
@@ 3.1.1 Neural Representations and SDF-Based Methods
Neural implicit representations, such as NeRF (Neural Radiance Fields) and NeuS (Neural Implicit Surfaces), have demonstrated state-of-the-art performance in view synthesis and geometric reconstruction. However, their uncertainty estimation techniques differ significantly.
@@
  • NeuS-based methods instead represent surfaces using a SDF, which provides a continuous and differentiable description of 3D geometry. However, existing NeRF uncertainty estimation techniques do not translate well to NeuS because they rely on ray-space density variations rather than surface-level geometric uncertainty.
@@ 3.2 Uncertainty Estimation in Ensemble-Based Learning
  • Computational Cost: Training multiple neural implicit models is memory- and computeintensive, particularly for high-resolution 3D reconstructions. Traditional ensemble methods require storing and evaluating multiple models, leading to scalability issues.
@@
  • Ray vs. Volume-Based Estimation: Many existing ensemble-based uncertainty estimation techniques in NeRF-like models operate strictly along rays. This results in uncertainty that does not align well with geometric structures, making it less interpretable.
While leveraging the strengths of ensemble-based learning and addressing key limitations in 3D implicit representations, BayesSDF establishes a new paradigm for uncertainty quantification in neural implicit fields.
@@ 3.3 Bayesian Inference for Ray-Based Sampling
BayesRays [2] is a recent approach for uncertainty quantification in neural radiance fields. It modifies NeRF by introducing a Bayesian framework that estimates uncertainty in the color and density outputs of the model. While BayesRays is effective for quantifying NeRF uncertainty, it remains tied to volumetric density representations and does not explicitly model surface uncertainty. Additionally, BayesRays relies on ray-space sparsification techniques, which can lead to misalignment between estimated uncertainty and actual geometric uncertainty. Unlike BayesRays, BayesSDF integrates reconstruction methods that inherently reveal volumetric qualities, ensuring direct alignment between uncertainty estimates and the scene geometry. This approach places less emphasis on rendering estimations, making it inherently more interpretable for geometric reconstruction tasks by exploiting surface variability rather than depth-based uncertainty.
@@
Table 1: Comparison of existing 3D reconstruction methods, evaluating their representation, uncertainty estimation, computational cost, and geometric consistency.
@@ 5.1 Theory
Our goal is to model an associated uncertainty measure, derived from the local sensitivity of the rendering function to small perturbations in the underlying deformation parameters. Let $\mathbf{x}\in\mathbb{R}^{3}$ denote a 3D point in scene space, and let $\mathbf{c}(\mathbf{x})\in\mathbb{R}^{3}$ be the color (in Red Green Blue (RGB)) rendered by the neural field. Rather than relying purely on a deterministic mapping $\mathbf{x}\mapsto\mathbf{c}(\mathbf{x})$, we attach a set of local deformation parameters $\mathbf{d}(\mathbf{x})$ that influence the geometry (and thus the color) through a learned distortion or deformation field. Formally, we assume $\mathbf{c}(\mathbf{x})\;=\;\mathcal{R}\bigl{(}\mathbf{x},\mathbf{d}(\mathbf{x});\Theta\bigr{)}$ where $\mathcal{R}$ is the rendering operator that depends on both the scene geometry (parameterized by a neural SDF model and any additional shape parameters $\Theta)$ and the local deformation $\mathbf{d}(\mathbf{x})$. To estimate uncertainty, we measure how sensitively $\mathbf{c}(\mathbf{x})$ changes with respect to changes in $\mathbf{d}(\mathbf{x})$. Intuitively, if small variations in $\mathbf{d}(\mathbf{x})$ induce large changes in color, we interpret that region as having higher uncertainty in the final scene reconstruction.
5.2 Distortion Fields
@@
The distortion field $f_{\mathrm{hash}}(\mathbf{x})$ captures spatial distortions in the SDF field, serving as a basis for uncertainty estimation. To represent the deformation parameters $\mathbf{d}(\mathbf{x})$ on a continuous domain, we employ a HashEncoding that functions as a voxel-like grid implemented through a hash table $\mathbf{d}(\mathbf{x})=f_{\mathrm{hash}}(\mathbf{x})\in\mathbb{R}^{3}$ where $f_{\mathrm{hash}}$ is a function learned by a single-level hash encoding with: $\mathrm{min\_res}=\mathrm{max\_res}=2^{\ell}$ for level-of-detail $\ell$, a hash table size of $2^{3\ell+1}$ entries, linear interpolation between grid corners, and 3 output channels per queried grid vertex. Effectively, for a given point $\mathbf{x}$, the hash encoding looks up an 8-corner neighborhood around the voxel containing $\mathbf{x}$, interpolates among these corners (each storing a 3D deformation vector), and outputs a single 3D vector $\mathbf{d}(\mathbf{x})$. This approach is parameter-efficient while remaining flexible. We denote $\mathbf{aabb}\in\mathbb{R}^{6}$ as the scene bounding box, which is used to normalize $\mathbf{x}$ to a unit cube or appropriate domain before querying the hash grid. Specifically, we generate a ray bundle from the dataset’s rays, containing origins, directions, and near/far bounds. We then use Neus’ hierarchical or “proposal” sampler to refine sample locations along each ray, focusing more samples near scene geometry (i.e., near the SDF zero level set). Let $\{\mathbf{x}_{i}\}$ be the set of sampled 3D points on a ray. When each ray sample $\mathbf{x}\in\mathbb{R}^{3}$ is generated, we transform $\mathbf{x}$ into normalized coordinates $\tilde{\mathbf{x}}=\mathrm{normalize}(\mathbf{x},\mathbf{aabb})$. Then, $\mathbf{d}(\mathbf{x})\;=\;\sum_{c=1}^{8}\alpha_{c}(\tilde{\mathbf{x}})\,\mathbf{d}_{c}\quad\text{where}\quad\bigl{\{}\mathbf{d}c\bigr{\}}{c=1}^{8}\;$ are are the corner deformation vectors that are found from a grid cell index in each dimension (one for each binary combination of $\lfloor\cdot\rfloor or\lceil\cdot\rceil$ in (x,y,z)) and $\alpha_{c}(\tilde{\mathbf{x}})$ are the trilinear interpolation coefficients $(\alpha_{c}(\mathbf{x}i)\;\in\;[0,1],\quad c=1,\ldots,8,)$ such that $\sum{c=1}^{8}\alpha_{c}(\mathbf{x}_{i})=1$. These coefficients are used to blend the deformation vectors from the 8 corners. This procedure is vital not only for generating $\mathbf{d}(\mathbf{x}_{i})$, but also for accumulating partial derivatives. Because each point’s deformation influences eight corner entries, we track which corners are involved and in what proportion.